OpenStudy (anonymous):

I need help! D: my problem is: what is the value of x1 in the equation y-6=1/2(x+2)

OpenStudy (a_clan):

-2

OpenStudy (anonymous):

can you please explain?

OpenStudy (a_clan):

compare the above equation, with standard form , (y - y1) = m (x -x1)

OpenStudy (anonymous):

:D THANK YOU SO SO MUCH!(:

OpenStudy (anonymous):

If what you mean is \[y-6=1/(2(x+2))\] Multiply by (x+2) and divide by (y-6) on both sides \[(y-6)(x+2)/(y-6)=(x+2)/(2(x+2)(y-6))\] Cancel out what needs to be cancelled out: \[(x+2)=1/(2(y-6))\] Subtract 2 from both sides \[x+2-2=(1/(2(y-6)))-2\] \[x=(1/(2(y-6)))-2\]